# The Bayesian Geometry of Transformer Attention

### v2

@@ Abstract.
Transformers often appear to perform Bayesian reasoning in context, but verifying this rigorously has been impossible: natural data lack analytic posteriors, and large models conflate reasoning with memorization. We address this by constructing Bayesian wind tunnels—controlled environments where the true posterior is known in closed form and memorization is provably impossible. In these settings, small transformers reproduce Bayesian posteriors with $10^{-3}$$10^{-4}$ bit accuracy, while capacity-matched MLPs fail by orders of magnitude, establishing a clear architectural separation.
Across two tasks—bijection elimination and Hidden Markov Model (HMM) state tracking—we find that transformers implement Bayesian inference through a consistent geometric mechanism: residual streams serve as the belief substrate, feed-forward networks perform the posterior update, and attention provides content-addressable routing. Geometric diagnostics reveal orthogonal key bases, progressive query–key alignment, and a low-dimensional value manifold parameterized by posterior entropy. During training this manifold unfurls while attention patterns remain stable, a frame–precision dissociation predicted by recent gradient analyses.
Taken together, these results demonstrate that hierarchical attention realizes Bayesian inference by geometric design, explaining both the necessity of attention and the failure of flat architectures. Bayesian wind tunnels provide a foundation for mechanistically connecting small, verifiable systems to reasoning phenomena observed in large language models.
@@ 1. Introduction
Do transformers perform Bayesian inference, or do they merely imitate it through pattern matching? Natural language offers no ground-truth posterior against which to verify predictions, and modern LLMs are too large and too entangled with their data to separate genuine probabilistic computation from memorization. Even when models behave Bayesianly, there is no direct way to confirm that the internal computation matches Bayes’ rule.
@@
    the hypothesis space is so large that memorization is impossible,
@@
Two wind tunnels. We study two settings of increasing difficulty:
@@
Transformers achieve machine-level Bayesian consistency in both. Capacity-matched MLPs trained identically fail catastrophically in both.
Mechanistic discovery. Across tasks, transformers implement Bayesian inference through a unified three-component architecture:
@@
    Residual stream as belief state: posterior information accumulates layer-by-layer.
@@
    Feed-forward networks as Bayesian update: FFNs perform the numerical posterior computation.
@@
    Attention as routing: QK geometry retrieves the relevant components of the belief for each update.
Geometric diagnostics reveal orthogonal key axes, progressive query–key alignment, and a one-dimensional value manifold that unfolds during training. These observations match predictions from recent gradient-based analyses of transformer learning.
Contribution. This paper provides the first empirical proof that transformers can realize exact Bayesian posteriors, identifies the geometric mechanism by which this occurs, and introduces Bayesian wind tunnels as a tool for probing algorithmic reasoning in small, verifiable settings.
Clarification on “Bayesian inference.” We do not claim a Bayesian posterior over network weights; we show that the learned predictor implements the Bayesian posterior predictive over latent task variables—the filtering posterior over hidden states in HMMs, or the elimination posterior over bijections. This is a statement about the input-output function the transformer computes, not about weight-space uncertainty.
@@ 2.3. Application to the bijection wind tunnel
In the bijection task, each $\theta$ is a bijection $\pi:\{1,\dots,V\}\to\{1,\dots,V\}$. A training sequence reveals k1k-1 input—-output pairs. Let $\mathcal{O}_{k-1}$ be the set of outputs already observed. Because each input appears at most once per sequence, the current query $x_{k}$ has never been seen before, so Bayes’ rule reduces to:
@@
producing a monotone staircase that shrinks by one bit whenever a mapping is revealed.
This closed-form posterior allows direct, position-by-position comparison between model entropy and Bayesian entropy; memorization is impossible because the hypothesis space size $V!$ is enormous.
@@ 2.4. Application to the HMM wind tunnel
The analytic posterior entropy is therefore
(7) $H_{\text{Bayes}}(t)~=~-\sum_{s=1}^{S}\alpha_{t}(s)\log_{2}\alpha_{t}(s).$
Because every training sequence is generated from a freshly sampled $(T,E)$, the hypothesis space is massive and memorization is impossible. The model must learn to (i) parse the header encoding $T$ and $E$, and (ii) implement a recursive Bayesian update.
@@ 2.5. Implications for model evaluation
The theoretical results above imply a practical diagnostic: a model that achieves the correct posterior entropy at every position is functionally Bayesian—it produces predictions with the same uncertainty profile as the exact posterior. [^1] Entropy-matching is necessary but not sufficient for distributional convergence in general, since different distributions can share the same entropy. However, direct verification via KL divergence confirms that our trained transformers match the full Bayesian posterior: $\text{KL}(\text{model}\|\text{Bayes})\approx 0$ nats and total variation distance $3\%$ for new keys, with exact retrieval (KL =0=0, TVD =0=0) for repeated keys in the with-replacement setting. Combined with the cross-entropy training objective (whose unique population minimizer is the Bayesian posterior predictive), low entropy calibration error provides strong evidence for Bayesian computation.
@@
therefore provides a direct, bit-level measure of Bayesian correctness, independent of accuracy or perplexity.
In later sections we show that transformers achieve near-perfect calibration, while matched MLPs do not.
@@ 3. Experimental Design
We evaluate whether small transformers can realize exact Bayesian inference by placing them in two controlled “Bayesian wind tunnels” where memorization is impossible and the analytic posterior is known in closed form. The two tasks—-bijection learning and Hidden Markov Model (HMM) state tracking—- probe different inference structures. Bijections require discrete hypothesis elimination; HMMs require recursive integration of stochastic transitions and emission likelihoods.
Across both settings, the evaluation criterion is simple: does the model’s predictive entropy $H_{\text{model}}$ match the analytic posterior entropy $H_{\text{Bayes}}$ at every position?
We measure this using mean absolute entropy error (MAE),
@@ 3.2. Task 2: Hidden Markov Model State Tracking
The second wind tunnel probes a qualitatively different inferential structure: recursive belief updating. Each sequence is derived from a fresh HMM with $S=5$ hidden states and $V=5$ observation symbols. Transition rows and emission rows are drawn independently from a symmetric Dirichlet distribution with all concentration parameters equal to 1 (i.e., $\text{Dirichlet}(1,1,1,1,1)$), ensuring diverse and non-degenerate dynamics.
@@
  • $K$ observation—-prediction pairs, each consisting of:
@@ 3.3. Architectures
To isolate the role of attention, we train MLPs with:
  • 18—-20 layers,
  • width 384—-400,
  • residual connections and layer normalization,
  • identical embedding layers and training protocol as the transformers.
Parameter counts match transformers within 1%. These MLPs serve as controls testing whether hierarchical attention is essential for the task.
@@ 3.4. Training Protocol
AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay 0.01, gradient clipping at 1.0.
Learning rates.
  • Bijections: constant $10^{-3}$.
  • HMMs: $3\times 10^{-4}$ with 1000-step warmup and cosine decay.
@@ 4. Results: Transformers Track the Bayesian Posterior
We evaluate whether transformers lie on the analytic Bayesian manifold using two behavioral tests: (1) pointwise calibration—does $H_{\text{model}}(t)$ match $H_{\text{Bayes}}(t)$ at every position? (2) generalization—does the learned computation extend to unseen bijections, unseen HMMs, and longer sequences?
We present results for bijections and HMMs in parallel, followed by MLP controls and multi-seed robustness.
@@ 4.1. Bijection Wind Tunnel: Exact Hypothesis Elimination
overlaid on $H_{\text{Bayes}}(k)=\log_{2}(V-k+1)$. The curves coincide across all positions, including late steps where only 2–4 hypotheses remain.
Quantitatively, the transformer achieves
@@
averaged over 2,000 held-out bijections. This error is smaller than single-precision numerical noise in the analytic posterior.
@@ 4.2. HMM Wind Tunnel: Recursive Bayesian State Tracking
The two curves are visually indistinguishable (Figure 5).
@@
To test algorithmic generalization, we roll the model out to $1.5\times$ and $2.5\times$ training length. The transformer remains remarkably close to the analytic posterior:

### v3

@@ Abstract.
Modern sequence models often appear to behave as Bayesian learners, but it remains unclear whether this reflects genuine probabilistic inference or task-specific heuristics. We introduce Bayesian wind tunnels—controlled environments where the true posterior is known in closed form and memorization is provably impossible—to resolve this question empirically. In these settings, small transformers reproduce exact Bayesian posteriors for filtering and hypothesis elimination with $10^{-3}$$10^{-4}$ bit accuracy, while capacity-matched MLPs fail by orders of magnitude. To understand which architectural ingredients enable exact inference, we decompose Bayesian computation into three inference primitives: (i) belief accumulation—integrating evidence into a running posterior; (ii) belief transport—propagating beliefs forward through stochastic dynamics; and (iii) random-access binding—retrieving stored hypotheses by content rather than position. Different tasks demand different subsets of these primitives, and different architectures can realize different subsets. Comparing Transformers, Mamba, LSTMs, and MLPs across bijection learning, HMM filtering, and associative recall, we find that Transformers realize all three primitives; Mamba realizes accumulation and transport but struggles with random-access binding; LSTMs realize only accumulation (of static sufficient statistics); and MLPs realize none. Geometric diagnostics reveal orthogonal key bases, low-dimensional value manifolds parameterized by posterior entropy, and—in Mamba—five discrete clusters corresponding to HMM hidden states. These results demonstrate that Bayesian computation is not monolithic: its realizability depends on the inference primitives a task demands and the architectural mechanisms available to implement them. Bayesian wind tunnels provide a foundation for mechanistically connecting small, verifiable systems to reasoning phenomena observed in large language models.
@@ 1. Introduction
Can transformers perform exact Bayesian inference—filtering and hypothesis elimination—or do they merely approximate it through pattern matching? Natural language offers no ground-truth posterior against which to verify predictions, and modern LLMs are too large and too entangled with their data to separate genuine probabilistic computation from memorization. Even when models behave Bayesianly, there is no direct way to confirm that the internal computation matches Bayes’ rule.
@@
    the hypothesis space is so large that memorization is computationally infeasible,
@@
Four wind tunnels. We study four settings:
@@
  • Bayesian regression: a continuous inference problem with closed-form Gaussian posterior over linear weights.
  • Associative recall: a content-based retrieval task testing the binding primitive.
To understand which architectural ingredients enable Bayesian inference, we decompose it into three inference primitives:
@@
    Belief accumulation: integrating evidence into a running posterior (e.g., updating $P(\theta\mid x_{1:t})$ as observations arrive).
@@
    Belief transport: propagating beliefs forward through stochastic dynamics (e.g., HMM filtering where hidden states evolve).
@@
    Random-access binding: retrieving stored hypotheses by content rather than position (e.g., recalling a target given a probe cue).
Different tasks demand different subsets of these primitives, and different architectures can realize different subsets. We test four architectures: Transformers, Mamba (a selective state-space model), LSTMs, and MLPs. Figure 1 summarizes which architectures realize which primitives.
Figure 1. Inference primitives taxonomy. Each row is a primitive; each column is an architecture. Transformers realize all three primitives. Mamba realizes accumulation and transport but struggles with random-access binding. LSTMs realize only accumulation of static sufficient statistics. MLPs realize none.
Findings. Transformers realize all three primitives and succeed on all tasks. Mamba realizes accumulation and transport—achieving state-of-the-art on HMM filtering—but struggles with random-access binding in associative recall. LSTMs realize only accumulation of static sufficient statistics: they succeed on belief revision (where the statistic is fixed-dimensional) but fail when the sufficient statistic itself evolves under dynamics or must be indexed by content. MLPs realize none and fail uniformly. Geometric diagnostics reveal orthogonal key axes, progressive alignment, and a low-dimensional value manifold parameterized by entropy. On HMM tracking, Mamba’s final layer organizes into five clusters—one per hidden state—showing that the model discovers the corner geometry of the belief simplex.
Contribution. This paper provides the first empirical proof that neural sequence models can realize exact Bayesian posteriors, introduces Bayesian wind tunnels as a tool for probing algorithmic reasoning in verifiable settings, and establishes a taxonomy of inference primitives that explains which architectures succeed on which tasks. The primitives framework unifies prior observations: content-based routing enables accumulation and transport; attention’s random-access capability additionally enables binding.
Clarification on “Bayesian inference.” We do not claim a Bayesian posterior over network weights; we show that the learned predictor implements the Bayesian posterior predictive over latent task variables—the filtering posterior over hidden states in HMMs, or the elimination posterior over bijections. This is a statement about the input-output function the transformer computes, not about weight-space uncertainty. Specifically, we demonstrate exact Bayesian inference in tasks whose posteriors factorize sequentially over the input—filtering and elimination, not general Bayesian model selection.
Roadmap: this paper as Lemma 1. This paper establishes the existence and internal geometry of exact Bayesian inference in transformers under verifiable conditions. It is the first of three papers forming a unified argument: Paper II shows that this geometry arises generically from gradient dynamics under cross-entropy training, explaining why transformers learn Bayesian structure. Paper III shows how these primitives compose in partially observed settings closer to natural language. Together, the trilogy characterizes when, why, and how neural sequence models implement probabilistic reasoning.
Structural Theorem (Inference Primitives and Architectural Realizability). Bayesian inference in sequential prediction decomposes into three primitives: (1) belief accumulation over fixed latent hypotheses, (2) belief transport under latent state dynamics, and (3) random-access binding between beliefs and past observations. Neural sequence architectures differ not in whether they can approximate Bayesian inference, but in which primitives they can realize: Recurrent architectures (LSTMs) implement accumulation of static sufficient statistics, but fail when inference requires transport of probability mass under nontrivial dynamics. State-space models (Mamba) implement accumulation and transport, but fail when inference requires random-access binding to arbitrary past observations. Attention-based transformers implement all three primitives by externalizing belief as a geometric, addressable representation rather than compressing it into fixed-size state. The dominance of transformers in reasoning tasks arises not from scale alone, but from primitive completeness: they are the minimal architecture realizing the full set of inference primitives.
@@ 2.3. Application to the bijection wind tunnel
In the bijection task, each $\theta$ is a bijection $\pi:\{1,\dots,V\}\to\{1,\dots,V\}$. A training sequence reveals k1k-1 input–output pairs. Let $\mathcal{O}_{k-1}$ be the set of outputs already observed. Because each input appears at most once per sequence, the current query $x_{k}$ has never been seen before, so Bayes’ rule reduces to:
@@
producing a monotone staircase that shrinks by one bit whenever a mapping is revealed. This closed-form posterior allows direct, position-by-position comparison between model entropy and Bayesian entropy; memorization is computationally infeasible because the hypothesis space size $V!$ is enormous.
@@ 2.4. Application to the HMM wind tunnel
Since models predict the next observation $o_{t+1}$ rather than hidden states, we evaluate the predictive entropy:
(7) $H_{\text{Bayes}}(t)~=~-\sum_{o}p(o_{t+1}\mid o_{1:t})\log_{2}p(o_{t+1}\mid o_{1:t}),$
where $p(o_{t+1}\mid o_{1:t})=\sum_{s,s^{\prime}}\alpha_{t}(s)T(s^{\prime}\mid s)E(o_{t+1}\mid s^{\prime})$. Because every training sequence is generated from a freshly sampled $(T,E)$, the hypothesis space is massive and memorization is computationally infeasible. The model must learn to (i) parse the header encoding $T$ and $E$, and (ii) implement a recursive Bayesian update.
2.5. Application to the regression wind tunnel
In the regression task, each $\theta$ is a weight vector $w\in\mathbb{R}^{d}$ with prior $w\sim\mathcal{N}(0,I)$. Given context observations $(x_{i},y_{i})$ where $y_{i}=w^{\top}x_{i}+\epsilon_{i}$ with $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, the posterior over $w$ is Gaussian with closed-form mean and covariance. The posterior predictive at any query point $x$ is also Gaussian, enabling exact computation of the Bayesian predictive distribution for comparison.
2.6. Application to the associative recall wind tunnel
In the associative recall task, each sequence contains $N$ cue–target pairs $(c_{i},t_{i})$ followed by probe cues. Unlike the other tasks which test belief accumulation or transport, associative recall tests the binding primitive: the model must store all pairs and retrieve by content when the probe arrives. Success requires content-based routing—identifying which stored pair matches the probe—rather than sequential Bayesian updating.
@@ 2.7. Implications for model evaluation
The theoretical results above imply a practical diagnostic: a model that achieves the correct posterior entropy at every position is functionally Bayesian—it produces predictions with the same uncertainty profile as the exact posterior. Combined with the cross-entropy training objective (whose unique population minimizer is the Bayesian posterior predictive), low entropy calibration error provides strong evidence for Bayesian computation.
Beyond entropy: full distributional verification.
Entropy-matching is necessary but not sufficient for distributional convergence in general, since different distributions can share the same entropy. In our wind tunnels, however, entropy serves as a diagnostic sufficient statistic: the structured nature of the tasks (discrete elimination over bijections, recursive filtering over HMM states) strongly constrains the space of distributions achieving a given entropy. To confirm full distributional equivalence, we directly verify via KL divergence that our trained transformers match the complete Bayesian posterior (Table 1). Throughout this paper we report entropy MAE as our primary metric because it provides an interpretable, bit-level measure that generalizes across tasks.
Table 1. Full distributional verification: KL divergence (nats) and total variation distance between transformer predictions and exact Bayesian posteriors. Bijection results are stratified by key type: “new” keys (first occurrence) vs. “repeated” keys (seen before in the sequence). Values are averaged over 2,000 held-out sequences.
[…]
Bijection task
     New keys $8.2\times 10^{-4}$ 0.027
     Repeated keys $10^{-6} <$$10^{-4}$
HMM task ($K=20$)
     All positions $1.1\times 10^{-4}$ 0.018
@@
therefore provides a direct, bit-level measure of Bayesian correctness, independent of accuracy or perplexity. In later sections we show that transformers achieve near-perfect calibration, while matched MLPs do not.
@@ 3. Experimental Design
We evaluate whether small transformers can realize exact Bayesian inference by placing them in four controlled “Bayesian wind tunnels” where memorization is computationally infeasible and the analytic posterior is known in closed form. The four tasks—bijection learning, Hidden Markov Model (HMM) state tracking, Bayesian regression, and associative recall— probe different inference structures. Bijections require discrete hypothesis elimination; HMMs require recursive integration of stochastic transitions and emission likelihoods. Across both settings, the evaluation criterion is simple: does the model’s predictive entropy $H_{\text{model}}$ match the analytic posterior entropy $H_{\text{Bayes}}$ at every position? We measure this using mean absolute entropy error (MAE),
@@ 3.2. Task 2: Hidden Markov Model State Tracking
The second wind tunnel probes a qualitatively different inferential structure: recursive belief updating. Each sequence is derived from a fresh HMM with $S=5$ hidden states and $V=5$ observation symbols. Transition rows and emission rows are drawn independently from a symmetric Dirichlet distribution with all concentration parameters equal to 1 (i.e., $\text{Dirichlet}(1,1,1,1,1)$), ensuring diverse and non-degenerate dynamics. The initial state distribution $\pi_{0}$ is also drawn from $\text{Dirichlet}(1,1,1,1,1)$.
@@
  • $K$ observation–prediction pairs, each consisting of:
@@
3.3. Task 3: Associative Recall
To isolate the random-access binding primitive, we use a standard associative recall task: the model must retrieve a target given a probe cue from a set of cue–target pairs presented in context.
Task format.
Each sequence contains $N$ cue–target pairs $(c_{i},t_{i})$ followed by $Q$ probe cues. Cues and targets are drawn uniformly from disjoint vocabularies of size 256 each (total vocabulary 522 including special tokens). The model must predict the associated target for each probe. With $N=16$ pairs and $Q=3$ probes, sequence length is 137 tokens.
Experimental setup.
  • Training: 50,000 sequences, 30 epochs
  • Models: Transformer, Mamba, LSTM with $d_{\mathrm{model}}=128$, comparable parameter counts ($\sim$570k–660k)
  • Metric: Exact-match accuracy on probe predictions
Why this tests binding.
Unlike bijection (where the hypothesis space shrinks predictably) or HMM (where beliefs evolve through dynamics), associative recall requires late, content-dependent retrieval: the model cannot know which cue–target pairs matter until the probe arrives. This requires storing all pairs and retrieving by content—the binding primitive.
Results preview.
[…]
3.4. Task 4: Bayesian Regression
To test continuous latent variables, we use multivariate linear regression with a Gaussian prior.
Task format.
Each sequence derives from a fresh linear regression with weights $w\sim\mathcal{N}(0,I_{d})$ where $d=3$. Observations follow $y_{i}=w^{\top}x_{i}+\epsilon_{i}$ with $\epsilon_{i}\sim\mathcal{N}(0,0.25)$ and inputs $x_{i}\sim\mathcal{N}(0,I_{d})$. The model observes $k=6$ context pairs before predicting at a query point drawn from the same input distribution. Outputs are discretized into 41 bins uniformly spaced over [5,5][-5,5], covering $99.9\%$ of the predictive mass given the prior and noise scales.
Bayesian ground truth.
The posterior over $w$ given context is Gaussian with closed-form mean and covariance, yielding a Gaussian predictive distribution at any query point. This enables exact computation of the Bayesian predictive for comparison.
Why this tests continuous inference.
Unlike bijection (discrete hypotheses) or HMM (discrete states), regression requires inference over a continuous parameter space. The Gaussian structure ensures the posterior remains tractable while testing whether transformers can calibrate uncertainty in continuous settings.
@@ 3.5. Architectures
Mamba (selective state-space model).
To test whether attention specifically is required, or whether content-based routing more generally suffices, we train Mamba models (Gu and Dao, 2023):
  • Bijection Mamba (3.77M): 9 layers, $d_{\mathrm{model}}=256$, state dimension 16.
  • HMM Mamba (3.77M): 9 layers, $d_{\mathrm{model}}=256$, state dimension 16.
Mamba replaces attention with a selective state-space mechanism: input-dependent matrices $(\Delta,B,C)$ gate the recurrent state update. This provides content-based routing without explicit query-key matching.
LSTM baselines.
To test whether recurrence alone suffices, we train LSTMs:
  • Bijection LSTM (4.77M): 9 layers, hidden dimension 256.
  • HMM LSTM (4.77M): 9 layers, hidden dimension 256.
LSTMs have recurrent state but use fixed gating—forget, input, and output gates do not depend on content relationships across sequence positions. This tests whether recurrence without content-based routing can implement Bayesian inference.
@@
To isolate the role of sequence structure entirely, we train MLPs with 18–20 layers, width 384–400, residual connections and layer normalization (parameter counts match transformers within 1%). The MLP receives the entire context sequence (all previous tokens concatenated) as input at each prediction position, but processes this flattened input without any attention or recurrence—testing whether feedforward computation over concatenated context suffices for Bayesian inference. We chose this design to give the MLP maximal access to context information; stronger attention-free baselines (e.g., Perceiver-style pooling, permutation-invariant aggregation) could narrow the gap but would reintroduce forms of content-based routing.
@@ 3.6. Training Protocol
AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay 0.01, gradient clipping at 1.0. Batch size is 64 for all tasks.
Learning rates and training steps.
  • Bijections: constant $10^{-3}$ for 150k steps.
  • HMMs: $3\times 10^{-4}$ with 1000-step warmup and cosine decay for 100k steps.
  • Regression: $5\times 10^{-4}$ for 50k steps.
@@ 4. Results: Transformers Track the Bayesian Posterior
We evaluate whether transformers lie on the analytic Bayesian manifold using two behavioral tests: (1) pointwise calibration—does $H_{\text{model}}(t)$ match $H_{\text{Bayes}}(t)$ at every position? (2) generalization—does the learned computation extend to unseen bijections, unseen HMMs, and longer sequences? We present results for bijections and HMMs in parallel, followed by MLP controls and multi-seed robustness.
@@ 4.1. Bijection Wind Tunnel: Exact Hypothesis Elimination
overlaid on $H_{\text{Bayes}}(k)=\log_{2}(V-k+1)$. The curves coincide across all positions, including late steps where only 2–4 hypotheses remain. Quantitatively, the transformer achieves
@@
averaged over 2,000 held-out bijections. This error is smaller than single-precision numerical noise in the analytic posterior. All analytic Bayes computations use double-precision arithmetic, so reported errors are meaningful. We verify full distributional agreement via KL and TVD: across all positions, $\text{KL}(\text{model}\|\text{Bayes})<0.01$ nats and TVD $3\%$, with per-position analysis confirming agreement holds across the full entropy range (high, moderate, and low entropy positions).
@@ 4.2. HMM Wind Tunnel: Recursive Bayesian State Tracking
The two curves are visually indistinguishable (Figure 7). (Note: The architecture comparison in Table 2 uses a smaller parameter-matched model at 2.7M for fair cross-architecture comparison, yielding MAE =0.049=0.049 bits; the result here uses the task-optimized 2.68M transformer.)
@@
To test algorithmic generalization, we roll the model out to $1.5\times$ and $2.5\times$ training length. The transformer stays close to the analytic posterior: